OpenStudy (anonymous):

the exact value of sin13pi/8 is?

OpenStudy (sirm3d):

\[\frac{13\pi}{8}=\frac{1}{2}\left(\frac{13\pi}{4}\right)\]

OpenStudy (sirm3d):

you will need this identity \[\sin\left(\frac{1}{2}A\right)=\pm \sqrt{\frac{1-\cos A}{2}}\] choose the sign that corresponds to the the sine function on the quadrant where \(13\pi/8\) lies